TC=50+10Q+0.04Q^2
MC=10+0.08Q
TR=40+20Q+0.02Q^2
MR=20+0.04Q

1) Find the quantity that the firm produces

2)Find the price at which the firm produces

3)Find the total cost,total variable cost and total fixed cost

4) how much does the firm need to earn to avoid from closing its operations
1
Expert's answer
2019-10-22T09:38:23-0400

1) The firm produces such quantity, at which MR = MC, so:

10 + 0.08Q = 20 + 0.04Q,

0.04Q = 10,

Q = 250 units.

2) The price at which the firm produces is found from the demand curve.

MR = 20+0.04Q, so the demand equation is P = 20 + 0.02Q = 20 + 0.02×250 = 25.

3) The total cost is:

TC = 50+10Q+0.04Q^2 = 50 + 10×250 + 0.04×250^2 = 5050.

Total variable cost is TVC = 10Q+0.04Q^2 = 5000,

and total fixed cost is TFC = 50.

4) The firm need to produce at P > AVC to avoid from closing its operations.